# Learning Adaptive and View-Invariant Vision Transformer with Multi-Teacher Knowledge Distillation for Real-Time UAV Tracking

### v1

@@ Abstract
Visual tracking has made significant strides due to the adoption of transformer-based models. Most state-of-the-art trackers struggle to meet real-time processing demands on mobile platforms with constrained computing resources, particularly for real-time unmanned aerial vehicle (UAV) tracking. To achieve a better balance between performance and efficiency, we introduce AVTrack, an adaptive computation framework designed to selectively activate transformer blocks for real-time UAV tracking. The proposed Activation Module (AM) dynamically optimizes the ViT architecture by selectively engaging relevant components, thereby enhancing inference efficiency without significant compromise to tracking performance. Furthermore, to tackle the challenges posed by extreme changes in viewing angles often encountered in UAV tracking, the proposed method enhances ViTs’ effectiveness by learning view-invariant representations through mutual information (MI) maximization. Two effective design principles are proposed in the AVTrack. Building on it, we propose an improved tracker, dubbed AVTrack-MD, which introduces the novel MI maximization-based multi-teacher knowledge distillation (MD) framework. It harnesses the benefits of multiple teachers, specifically the off-the-shelf tracking models from the AVTrack, by integrating and refining their outputs, thereby guiding the learning process of the compact student network. Specifically, we maximize the MI between the softened feature representations from the multi-teacher models and the student model, leading to improved generalization and performance of the student model, particularly in noisy conditions. Extensive experiments on multiple UAV tracking benchmarks demonstrate that AVTrack-MD not only achieves performance comparable to the AVTrack baseline but also reduces model complexity, resulting in a significant 17% increase in average tracking speed. Codes is be available at https://github.com/wuyou3474/AVTrack.
Keywords: UAV tracking, real-time, vision transformer, activation module, view-invariant representations, multi-teacher knowledge distillation.
@@
1 Introduction
Unmanned Aerial Vehicle (UAV) tracking has become increasingly critical due to its diverse applications, including path planning (Lee and Hwang, 2015), public safety (Zhang et al., 2019), visual surveillance (Tang et al., 2017), environmental monitoring (Sharma et al., 2016), industrial inspections (Mourtzis et al., 2021), and agriculture (Singh and Sharma, 2022), etc. Specifically, UAV tracking involves determining and predicting the location of a specific object in successive aerial images, which are captured by mobile cameras at relatively high altitudes. As a result, common challenging scenarios have emerged in the UAV tracking community, including fast object or UAV motion, extreme viewing angles, motion blur, low resolution, and significant occlusions where objects are obscured by other elements. Moreover, the efficiency of tracking algorithms is crucial, as mobile platforms like UAVs have limited computational resources (Li et al., 2023; Wang et al., 2023; Ma et al., 2023; Liu et al., 2022b; Li et al., 2020). In order to meet the unique demands of UAV applications, a practical tracker should ensure accurate tracking while running with relatively low computational and power consumption.
Neither discriminative correlation filter (DCF)-based trackers nor deep learning (DL)-based trackers can achieve a satisfactory balance between performance and efficiency. DCF-based trackers offer higher efficiency but struggle with tracking accuracy and robustness, making them unsuitable for complex UAV tracking scenarios, while DL-based trackers are hindered by high computational costs and sluggish performance, limiting their practical applicability. Single-stream architecture has recently emerged as a popular strategy in DL-based trackers, seamlessly integrating feature extraction and fusion with pre-trained Vision Transformer (ViT) backbone networks. A number of representative methods, including OSTrack (Ye et al., 2022), SimTrack (Chen et al., 2022), Mixformer (Cui et al., 2022), DropMAE (Wu et al., 2023), ZoomTrack (Kou et al., 2023), AQATrack (Xie et al., 2024b), EVPTrack (Shi et al., 2024), and HIPTrack (Cai et al., 2024), demonstrate the immense potential of this paradigm shift in tracking task. Motivated by this, Aba-VTrack (Li et al., 2023) proposes an effective and efficient DL-based tracker for real-time UAV tracking based on this framework, utilizing an adaptive and background-aware token calculation technique to minimize inference time. However, the presence of unstructured access operations with a variable number of tokens results in significant time costs.
@@ 1 Introduction
On the other hand, we introduce a novel approach to learning view-invariant feature representations by maximizing the mutual information between the backbone features extracted from two different views of the target, enhancing the tracker’s robustness against viewpoint changes. Specifically, mutual information is a measure that quantifies the dependence or relationship between two variables (Steuer et al., 2002). Mutual information maximization involves enhancing the mutual information between different components or variables within a system and is widely used in various computer vision tasks (Liu et al., 2022c; Yang et al., 2022; Hjelm et al., 2018). However, to the best of our knowledge, the effectiveness of this strategy in UAV tracking has not been extensively explored. In our work, to obtain view-invariant representations, the tracker learns to preserve essential information about the target regardless of viewpoint changes by maximizing the mutual information between two different views of the target. By employing this method, we call the resulting representations view-invariant representations. In theory, models trained with these representations are better at generalizing across diverse viewing conditions, making them more robust in real-world scenarios with frequent viewpoint changes. View-invariant representations are useful to trackers for effective tracking since extreme changes in viewing angles are a common challenge in UAV tracking.
@@
To enhance the single-stream paradigm for real-time UAV tracking, we conduct extensive experiments, highlight the two key design principles mentioned above, and introduce an improved version called AVTrack-MD. Specifically, we introduce a novel multi-teacher knowledge distillation (MD) framework based on MI maximization in the AVTrack. In the proposed MD, three off-the-shelf tracking models from the AVTrack serve as teachers, distilling knowledge into a more lightweight student model. The student model maintains the same structure as the teachers but features a smaller ViT backbone, with half the number of ViT blocks. In practice, multiple teacher models offer comprehensive guidance by providing diverse knowledge, which is advantageous for training the student (Gou et al., 2021; Wang and Yoon, 2021). Multi-teacher knowledge distillation is an effective strategy that is widely used across various computer vision tasks, including image super-resolution (Jiang et al., 2024), image classification (Wen et al., 2024), and visual retrieval (Ma et al., 2024). However, to the best of our knowledge, its effectiveness in UAV tracking has not been extensively investigated. In our work, given the mean squared error (MSE) is sensitive to noise and outliers (Hinton et al., 2015), we maximize the MI between the aggregated softened features of the multi-teacher models and the student model’s softened features, leading to improved generalization and performance of the student model, particularly in noisy conditions. Extensive experimental results demonstrate that the proposed method can produce a high-performing student network, with overall tracking performance that is comparable to or even exceeds that of teacher networks, while using significantly fewer parameters and achieving faster speeds (See Fig. 1 and Table 1).
@@
This work extends our previous conference paper, i.e., AVTrack (Li et al., 2024b), accepted at ICML 2024. In this extended paper, we propose AVTrack-MD, an improved tracker that significantly enhances the efficiency of AVTrack while achieving the overall tracking performance comparable to or even superior to it. We also provide extensive experiments and detailed implementations. The new contributions of this extended work are summarized as follows:
@@
2 Preliminaries
We begin by reviewing the background of UAV tracking, view-invariant feature representation, and multi-teacher knowledge distillation tasks.
2.1 UAV Tracking
@@
There are two main types of modern UAV trackers: DCF-based and CNN-based trackers (Li et al., 2021; Liu et al., 2022a). DCF-based trackers are favored for UAV tracking due to their remarkable efficiency on CPU, but they face difficulties in maintaining robustness under complex conditions (Li et al., 2021, 2020; Huang et al., 2019). Current CNN-based trackers, like those in (Cao et al., 2021, 2022), show advancements in tracking precision and robustness for UAV tracking, but their efficiency lags behind DCF-based trackers considerably. Although model compression techniques, as seen in (Wang et al., 2022; Wu et al., 2022; Zhong et al., 2023), were utilized to enhance efficiency, these trackers still face challenges associated with unsatisfactory tracking precision. A recent trend in the visual tracking community shows an increasing preference for single-stream architectures that seamlessly integrate feature extraction and correlation using pre-trained ViT backbone networks (Xie et al., 2021; Cui et al., 2022; Ye et al., 2022; Xie et al., 2022, 2024b), demonstrating significant potential. Several representative methods, including OSTrack (Ye et al., 2022), Mixformer (Cui et al., 2022), DropMAE (Wu et al., 2023), ZoomTrack (Kou et al., 2023), AQATrack (Xie et al., 2024b), EVPTrack (Shi et al., 2024), etc., demonstrate the significant success of applying this paradigm to tracking tasks. Although these frameworks are efficient due to their compact nature, very few are based on lightweight ViTs, making them impractical for real-time UAV tracking. To address this, TATrack (Li et al., 2024a) proposes an efficient ViT-based tracking framework for real-time UAV tracking, which integrates feature learning and template-search coupling into an efficient one-stream ViT. Similarly, ETDMT (Zhang et al., 2024) builds on lightweight ViT to introduce a tracker that combines template distinction with temporal context, while Aba-ViTrack (Li et al., 2023) enhances efficiency in real-time UAV tracking using lightweight ViTs and an adaptive background-aware token computation method. However, the variable token number in (Li et al., 2023) necessitates unstructured access operations, leading to notable time costs. Recent research has focused on improving the efficiency of ViTs by balancing their representation capabilities with computational efficiency. Methods like lightweight ViTs, model compression, and hybrid designs (Wang et al., 2020; Zhang et al., 2022b; Mao et al., 2021; Chen et al., 2021b; Li et al., 2022c) have been explored, but they often sacrifice accuracy or require time-consuming fine-tuning. Recent developments in efficient ViTs with conditional computation focus on adaptive inference, dynamically adjusting computational load based on input complexity to accelerate model performance. For example, DynamicViT (Rao et al., 2021) introduces control gates to selectively process tokens, while A-ViT (Yin et al., 2022) employs an Adaptive Computation Time strategy to avoid auxiliary halting networks, achieving gains in efficiency, accuracy, and token prioritization. Aba-ViTrack (Li et al., 2023) effectively utilizes the latter, showcasing significant potential for real-time UAV tracking. In our work, we focus on improving the efficiency of ViTs for UAV tracking through more structured methods, specifically the adaptive activation of Transformer blocks for feature representation.
2.2 View-Invariant Feature Representation
@@
View-invariant feature representation has garnered significant attention in the field of computer vision and image processing. This technique aims to extract features from images or visual data that remain consistent across various viewpoints or orientations, providing robustness to changes in the camera angle or scene configuration (Kumie et al., 2024a; Bracci et al., 2018; Li et al., 2017; Rao et al., 2002). Traditional methods typically rely on handcrafted features and geometric transformations to achieve view-invariant representations (Xia et al., 2012a; Ji and Liu, 2010; Rao et al., 2002). While effective, these traditional methods are often limited to specific scenarios with relatively simple and fixed backgrounds, making them unsuited for handling the complexity and variability of real-world visual data. Recently, many researchers have been eager to apply Convolutional Neural Networks (CNNs) and other deep architectures for extracting view-invariant representations, thanks to deep learning’s remarkable ability to extract discriminative features (Kumie et al., 2024b; Men et al., 2023; Pak et al., 2023; Gao et al., 2022; Shiraga et al., 2016). For instance, Kumie et al. (2024b) propose the Dual-Attention Network (DANet) for view-invariant action recognition, incorporating relation-aware spatiotemporal self-attention and cross-attention modules to learn representative and discriminative action features. Gao et al. (2021) propose a View Transformation Network (VTN) that realizes the view normalization by transforming arbitrary-view action samples to a base view to seek a view-invariant representation. These approaches leverage the capacity of deep models to capture complex patterns and variations in visual data. By ensuring that learned features are resilient to changes in viewpoint, these methods contribute to the robustness and generalization of vision-based systems. View-invariant feature representation is helpful for a variety of vision tasks, such as action recognition (Xia et al., 2012b; Liu et al., 2018; Kumie et al., 2024a), pose estimation (Bracci et al., 2018), human re-identification (Liu et al., 2023; Perwaiz et al., 2023), and object detection (Feng et al., 2022). However, to the best of our knowledge, the integration of view-invariant representations into visual tracking frameworks remains underexplored. In our work, we make the first attempt to learn view-invariant feature representations through mutual information maximization based on ViTs, specifically tailored for UAV tracking. This marks the first instance where ViTs are employed to acquire view-invariant feature representations in the context of UAV tracking.
2.3 Multi-Teacher Knowledge Distillation
@@
Multi-teacher knowledge distillation (MD) leverages knowledge from multiple off-the-shelf pre-trained teacher models to guide the student, enhancing its generalization by inheriting diverse knowledge from various teachers (Gou et al., 2021; Wang and Yoon, 2021). It has primarily been studied in the contexts of image classification (You et al., 2017; Yuan et al., 2021; Lan et al., 2024; Wen et al., 2024), action recognition (Wu et al., 2019) and visual retrieval (Ma et al., 2024). For instance, Wen et al. (2024) proposed the multi-teacher distillation method for class incremental learning, using weight permutation, feature perturbation, and diversity regularization to ensure diverse mechanisms in teachers. To better leverage the complementary knowledge of each modality, Lan et al. (2024) propose a multi-teacher multi-modal knowledge distillation framework to guide the training of the multi-modal fusion network and further improve the multi-modal feature fusion process. Whiten-MTD (Ma et al., 2024) aligns the outputs of teacher models by whitening their similarity distributions and identifies the most effective fusion strategy for their multi-teacher distillation framework through empirical analysis. To the best of our knowledge, existing knowledge distillation methods for visual tracking rely on a single-teacher framework (Li et al., 2022a; Sun et al., 2023), limiting their ability to fully utilize the diverse knowledge from multiple teachers, potentially leading to suboptimal tracking performance. In this work, we propose a simple yet effective MI maximization-based multi-teacher knowledge distillation framework, integrated into AVTrack, to develop a more efficient UAV tracker. For the teacher models, we utilize all three off-the-shelf trackers from AVTrack (i.e., AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA), offering diverse and high-quality teachers without requiring additional operations, such as those in (Wen et al., 2024), to implement varied mechanisms in the teachers. To ensure the student model captures the most relevant information from the teacher models’ representations, we propose maximizing the MI between the averaged softened feature representation of the multi-teacher models and the student model’s softened feature representation. Averaging the predictions of all teacher models is a common and effective method, as it reduces biases and provides a more objective output than any single teacher’s prediction (Buciluǎ et al., 2006; Ba and Caruana, 2014; Ilichev et al., 2021; Song and Chai, 2018).

### v2

# Learning an Adaptive and View-Invariant Vision Transformer for Real-Time UAV Tracking

@@ Abstract
Visual tracking has made significant strides due to the adoption of transformer-based models. Most state-of-the-art trackers struggle to meet real-time processing demands on mobile platforms with constrained computing resources, particularly for real-time unmanned aerial vehicle (UAV) tracking. To achieve a better balance between performance and efficiency, we introduce AVTrack, an adaptive computation framework designed to selectively activate transformer blocks for real-time UAV tracking. The proposed Activation Module (AM) dynamically optimizes the ViT architecture by selectively engaging relevant components, thereby enhancing inference efficiency without significant compromise to tracking performance. Furthermore, to tackle the challenges posed by extreme changes in viewing angles often encountered in UAV tracking, the proposed method enhances ViTs’ effectiveness by learning view-invariant representations through mutual information (MI) maximization. Two effective design principles are proposed in the AVTrack. Building on it, we propose an improved tracker, dubbed AVTrack-MD, which introduces the novel MI maximization-based multi-teacher knowledge distillation (MD) framework. It harnesses the benefits of multiple teachers, specifically the off-the-shelf tracking models from the AVTrack, by integrating and refining their outputs, thereby guiding the learning process of the compact student network. Specifically, we maximize the MI between the softened feature representations from the multi-teacher models and the student model, leading to improved generalization and performance of the student model, particularly in noisy conditions. This improvement provides a better trade-off between performance and efficiency. Extensive experiments on multiple UAV tracking benchmarks demonstrate that AVTrack-MD not only achieves performance comparable to the AVTrack but also reduces model complexity, resulting in a significant 17% increase in average tracking speed. Codes is be available at https://github.com/wuyou3474/AVTrack.
Index Terms:
UAV tracking, real-time, vision transformer, activation module, view-invariant representations, multi-teacher knowledge distillation.
@@
I Introduction
Unmanned aerial vehicle (UAV) tracking has become increasingly critical due to its diverse applications, including path planning [1], public safety [2], visual surveillance [3], environmental monitoring [4], industrial inspections [5], and agriculture [6], etc. Specifically, UAV tracking involves determining and predicting the location of a specific object in successive aerial images, which are captured by mobile cameras at relatively high altitudes. As a result, common challenging scenarios have emerged in the UAV tracking community, including fast object or UAV motion, extreme viewing angles, motion blur, low resolution, and significant occlusions where objects are obscured by other elements. Moreover, the efficiency of tracking algorithms is crucial, as mobile platforms like UAVs have limited computational resources [7, 8, 9, 10, 11]. In order to meet the unique demands of UAV applications, a practical tracker should ensure accurate tracking while running with relatively low computational and power consumption.
Neither discriminative correlation filter (DCF)-based trackers nor deep learning (DL)-based trackers can achieve a satisfactory balance between performance and efficiency. DCF-based trackers offer higher efficiency but struggle with tracking accuracy and robustness, making them unsuitable for complex UAV tracking scenarios, while DL-based trackers are hindered by high computational costs and sluggish performance, limiting their practical applicability. Single-stream architecture has recently emerged as a popular strategy in DL-based trackers, seamlessly integrating feature extraction and fusion with pre-trained Vision Transformer (ViT) backbone networks. A number of representative methods, including OSTrack [12], SimTrack [13], Mixformer [14], DropMAE [15], ZoomTrack [16], AQATrack [17], EVPTrack [18], and HIPTrack [19], demonstrate the immense potential of this paradigm shift in tracking task. Motivated by this, Aba-VTrack [7] proposes an effective and efficient DL-based tracker for real-time UAV tracking based on this framework, utilizing an adaptive and background-aware token calculation technique to minimize inference time. However, the presence of unstructured access operations with a variable number of tokens results in significant time costs.
@@ I Introduction
On the other hand, we introduce a novel approach to learning view-invariant feature representations by maximizing the mutual information between the backbone features extracted from two different views of the target, enhancing the tracker’s robustness against viewpoint changes. Specifically, mutual information is a measure that quantifies the dependence or relationship between two variables [20]. Mutual information maximization involves enhancing the mutual information between different components or variables within a system and is widely used in various computer vision tasks [21, 22, 23]. However, to the best of our knowledge, the effectiveness of this strategy in UAV tracking has not been extensively explored. In our work, to obtain view-invariant representations, the tracker learns to preserve essential information about the target regardless of viewpoint changes by maximizing the mutual information between two different views of the target. By employing this method, we call the resulting representations view-invariant representations. In theory, models trained with these representations are better at generalizing across diverse viewing conditions, making them more robust in real-world scenarios with frequent viewpoint changes. View-invariant representations are useful to trackers for effective tracking since extreme changes in viewing angles are a common challenge in UAV tracking.
@@
To enhance the single-stream paradigm for real-time UAV tracking, we conduct extensive experiments, highlight the two key design principles mentioned above, and introduce an improved version called AVTrack-MD. Specifically, we introduce a novel multi-teacher knowledge distillation (MD) framework based on MI maximization in the AVTrack. In the proposed MD, three off-the-shelf tracking models from the AVTrack serve as teachers, distilling knowledge into a more lightweight student model. The student model maintains the same structure as the teachers but features a smaller ViT backbone, with half the number of ViT blocks. In practice, multiple teacher models offer comprehensive guidance by providing diverse knowledge, which is advantageous for training the student [24, 25]. Multi-teacher knowledge distillation is an effective strategy that is widely used across various computer vision tasks, including image super-resolution [26], image classification [27], and visual retrieval [28]. However, to the best of our knowledge, its effectiveness in UAV tracking has not been extensively investigated. In our work, given the mean squared error (MSE) is sensitive to noise and outliers [29], we maximize the MI between the aggregated softened features of the multi-teacher models and the student model’s softened features, leading to improved generalization and performance of the student model, particularly in noisy conditions. Extensive experimental results demonstrate that the proposed method can produce a high-performing student network, with overall tracking performance that is comparable to or even exceeds that of teacher networks, while using significantly fewer parameters and achieving faster speeds (See Fig. 1 and Table I).
@@
This work extends our previous conference paper, i.e., AVTrack [30], accepted at ICML 2024. In this extended paper, we propose AVTrack-MD, an improved tracker that significantly enhances the efficiency of AVTrack while achieving the overall tracking performance comparable to or even superior to it. We also provide extensive experiments and detailed implementations. The new contributions of this extended work are summarized as follows:
@@
II RELATED WORKS
II-A UAV Tracking
@@
There are two main types of modern UAV trackers: DCF-based and CNN-based trackers [31, 32]. DCF-based trackers are favored for UAV tracking due to their remarkable efficiency on CPU, but they face difficulties in maintaining robustness under complex conditions [31, 11, 33]. Current CNN-based trackers, like those in [34, 35], show advancements in tracking precision and robustness for UAV tracking, but their efficiency lags behind DCF-based trackers considerably. Although model compression techniques, as seen in [36, 37, 38], were utilized to enhance efficiency, these trackers still face challenges associated with unsatisfactory tracking precision. A recent trend in the visual tracking community shows an increasing preference for single-stream architectures that seamlessly integrate feature extraction and correlation using pre-trained ViT backbone networks [39, 14, 12, 40, 17], demonstrating significant potential. Several representative methods, including OSTrack [12], Mixformer [14], DropMAE [15], ZoomTrack [16], AQATrack [17], EVPTrack [18], etc., demonstrate the significant success of applying this paradigm to tracking tasks. Although these frameworks are efficient due to their compact nature, very few are based on lightweight ViTs, making them impractical for real-time UAV tracking. To address this, several studies have started exploring the use of efficient ViTs built on one-stream framework for real-time UAV tracking [7, 41, 42, 43]. Specifically, ETDMT [43] builds on lightweight ViT to introduce a tracker that combines template distinction with temporal context, while Aba-ViTrack [7] enhances efficiency in real-time UAV tracking using lightweight ViTs and an adaptive background-aware token computation method. However, the variable token number in [7] necessitates unstructured access operations, leading to notable time costs. Recent research has focused on improving the efficiency of ViTs by balancing their representation capabilities with computational efficiency. Methods like lightweight ViTs, model compression, and hybrid designs [44, 45, 46, 47, 48] have been explored, but they often sacrifice accuracy or require time-consuming fine-tuning. Recent developments in efficient ViTs with conditional computation focus on adaptive inference, dynamically adjusting computational load based on input complexity to accelerate model performance. For example, DynamicViT [49] introduces control gates to selectively process tokens, while A-ViT [50] employs an Adaptive Computation Time strategy to avoid auxiliary halting networks, achieving gains in efficiency, accuracy, and token prioritization. Aba-ViTrack [7] effectively utilizes the latter, showcasing significant potential for real-time UAV tracking. In our work, we focus on improving the efficiency of ViTs for UAV tracking through more structured methods, specifically the adaptive activation of Transformer blocks for feature representation.
II-B View-Invariant Feature Representation
@@
View-invariant feature representation has garnered significant attention in the field of computer vision and image processing. This technique aims to extract features from images or visual data that remain consistent across various viewpoints or orientations, providing robustness to changes in the camera angle or scene configuration [51, 52, 53, 54]. Traditional methods typically rely on handcrafted features and geometric transformations to achieve view-invariant representations [55, 56, 54]. While effective, these traditional methods are often limited to specific scenarios with relatively simple and fixed backgrounds, making them unsuited for handling the complexity and variability of real-world visual data. Recently, many researchers have been eager to apply Convolutional Neural Networks (CNNs) and other deep architectures for extracting view-invariant representations, thanks to deep learning’s remarkable ability to extract discriminative features [57, 58, 59, 60, 61]. For instance, [57] propose the Dual-Attention Network (DANet) for view-invariant action recognition, incorporating relation-aware spatiotemporal self-attention and cross-attention modules to learn representative and discriminative action features. [62] propose a View Transformation Network (VTN) that realizes the view normalization by transforming arbitrary-view action samples to a base view to seek a view-invariant representation. These approaches leverage the capacity of deep models to capture complex patterns and variations in visual data. By ensuring that learned features are resilient to changes in viewpoint, these methods contribute to the robustness and generalization of vision-based systems. View-invariant feature representation is helpful for a variety of vision tasks, such as action recognition [63, 64, 51], pose estimation [52], human re-identification [65, 66], and object detection [67]. However, to the best of our knowledge, the integration of view-invariant representations into visual tracking frameworks remains underexplored. In our work, we make the first attempt to learn view-invariant feature representations through mutual information maximization based on ViTs, specifically tailored for UAV tracking. This marks the first instance where ViTs are employed to acquire view-invariant feature representations in the context of UAV tracking.
II-C Multi-Teacher Knowledge Distillation
@@
Multi-teacher knowledge distillation (MD) leverages knowledge from multiple off-the-shelf pre-trained teacher models to guide the student, enhancing its generalization by inheriting diverse knowledge from various teachers [24, 25]. It has primarily been studied in the contexts of image classification [68, 69, 70, 27], action recognition [71] and visual retrieval [28]. For instance, [27] proposed the multi-teacher distillation method for class incremental learning, using weight permutation, feature perturbation, and diversity regularization to ensure diverse mechanisms in teachers. To better leverage the complementary knowledge of each modality, [70] propose a multi-teacher multi-modal knowledge distillation framework to guide the training of the multi-modal fusion network and further improve the multi-modal feature fusion process. Whiten-MTD [28] aligns the outputs of teacher models by whitening their similarity distributions and identifies the most effective fusion strategy for their multi-teacher distillation framework through empirical analysis. To the best of our knowledge, existing knowledge distillation methods for visual tracking rely on a single-teacher framework [72, 73], limiting their ability to fully utilize the diverse knowledge from multiple teachers, potentially leading to suboptimal tracking performance. In this work, we propose a simple yet effective MI maximization-based multi-teacher knowledge distillation framework, integrated into AVTrack, to develop a more efficient UAV tracker. For the teacher models, we utilize all three off-the-shelf trackers from AVTrack (i.e., AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA), offering diverse and high-quality teachers without requiring additional operations, such as those in [27], to implement varied mechanisms in the teachers. To ensure the student model captures the most relevant information from the teacher models’ representations, we propose maximizing the MI between the averaged softened feature representation of the multi-teacher models and the student model’s softened feature representation. Averaging the predictions of all teacher models is a common and effective method, as it reduces biases and provides a more objective output than any single teacher’s prediction [74, 75, 76, 77].